# AcousTools: A ‘Full-Stack’, Python-Based, Acoustic Holography Library

### v3

@@ IV.1 Setup
from acoustools.Mesh import mesh_to_board
@@
    = mesh_to_board(’./data/Sphere-lam1.stl’)
This code loads the mesh and collects the positions and normals. In addition to custom transducer arrays, Acoustools has three common transducer configurations pre-configured, all based on arrays of 256 transducers arranged into 16x16 grids. This includes a single array positioned below the working volume and directed up AcousticElements, a single array positioned above the working volume and pointed down HighSpeedBEM, and one with the combination of these two arrays pointing towards one another GSPAT. These are supported as pre-constructed variables without needing to load a mesh, for example:
    from acoustools.Utilities import
                                TOP_BOARD
@@
Figure 3: Examples of using different meshes as the definition for transducer arrays creating a focal point in the origin of the working area (high pressure region at the centre of each subfigure). For a given mesh, a transducer is placed at each cell centre with the normal of the mesh defining the transducer normal, with normals pointing into the working volume inside the mesh. a) flat array similar to a common setup used in practice AcousticElements, b) Sphere c) Cube d) Teapot. For each array, the mesh is d so that the limits of the mesh in the x-axis is $\pm$11.8cm, centred at the origin (except the flat array which is translated down in order to give a working volume above it) and the normals point towards the origin.
After the boards are defined, points at which sound should be focused. To define a single point, $N=1$, can be simply determined as
from acoustools.Utilities import
                            create_points
p = create_points(N=1,x=target_x,
                      y=target_y,
                      z=target_z)
@@
IV.2 Propagators
As discussed above, the two main propagators used for acoustic holography are the piston model and the boundary element method. The piston model, as described in Eq. 1, uses a free-space model and is implemented in AcousTools as
from acoustools.Utilities import
            forward_model_batched
F = forward_model_batched(p,board,
        p_ref=p_ref,norms=norms,k=k)
@@ IV.2 Propagators
from acoustools.Mesh import load_scatterer
from acoustools.BEM import compute_E
@@
Because these propagators share the same structure - both are matrices where the complex pressure at a point is obtained by the matrix-vector product of the propagator and the transducer activations - these can be swapped without requiring additional changes to the rest of the code. This allows a user to compare the two contexts with minimal changes to their code. For some solver (e.g., IBIB), the propagator is the only component that needs to be interchanged between the piston model and BEM contexts.
@@
IV.3 Solvers
AcousTools provides a number of solvers, giving further flexibility depending on the context of an application. High-speed solvers, such as multipoint naive and GS-PAT, offer quick computationGSPAT, while slower solvers, such as WGS, can generate focal points with small variance of pressure TGS. As discussed, the choice of which solver to use is a critical stage in the development process, and a wide variety that can be easily compared enables a developer to make an informed choice.
@@ IV.3 Solvers
from acoustools.Solvers import
    iterative_backpropagation, gspat, wgs
@@
AcousTools can leverage implicit-gradient descent algorithms to optimise more complex objectives Diff-PAT; Diff-PAT-Expanded, which can go beyond what projective algorithms can do. Implicit gradient-descent methods can minimise Gor’kov potential directly, produce fields with specific metrics at specific points, and even combine multiple objectives simultaneously. Although these methods are often slower than others Diff-PAT and may not be suitable for all cases, their ability to use custom objectives provides greater flexibility for the developer.
@@
Then this can be sent to the implicit-gradient descent solver to be optimised, the exact optimiser and parameters can be customised such as the optimiser, the learning rate, weights of each term in the objective, and learning rate schedulersCLR. These parameters can then be passed to the solver:
@@
        objective_params={’lambda’:1.32e10},
@@
This results in the sound field shown in Fig. 4, using the Adam optimiserAdam, a learning rate of $0.01$, 10001000 iterations and $\lambda=1.32\times 10^{10}$ which ensures each component has a similar magnitude. One point has a minima in the Gor’kov potential corresponding to a low pressure region sandwiched between two high pressure regions and the other point has a maxima of Gor’kov potential corresponding to a region of high pressure, as defined in the objective. This could be useful for multi-modal applications where a focal point is used for haptic applications while the minima of Gor’kov potential corresponds to a levitation trap that can hold a small bead AcousticElements.
Projective solvers, such as IBP and GS-PAT, could not fully create the field in this situation, as they produce high pressure regions at their targets. These can be converted to traps through levitation signatures AcousticElements, but such signatures convert all focal points instead of only one. As seen in Fig. 4, a focal point has a local minima just below the high pressure regions, and so a similar result to Fig. 4 could be produced by focusing slightly above the target trap position. However, using this method may lead to both sub-optimal solutions with smaller forces and is harder to customise the relative strengths of the trap and focus. Therefore, the implicit-gradient solver is an additional tool for a developer, and combining this with the other methods provides a full toolbox to use.
@@ IV.4 Analysis
pressure = propagate_abs(transducer_phases,
                p, board=board)
U = gorkov(transducer_phases, p, board=board)
@@
pressure = propagate(transducer_phases,
                p, board=board)
Computing Gor’kov potential here uses analytical gradients of the piston model instead of using finite differences and makes it significantly faster compared to the standard method using finite differences AcousticElements; HighSpeedBEM. The speed-up between the two methods is shown in Tab.2. A similar analytical solution is provided using the analytical gradient of BEM. For both propagators, an analytical method for computing the gradient of Gor’kov potential is also provided for computing force. A summary of these metrics can be seen in Fig. 5, showing that all these metrics from pressure to stiffness. In this case, the propagator used is the piston model and metrics use the analytical gradients where possible, similar visulisations could be created using BEM by simply swapping the propagators as discussed above.
For a developer, the ability to select the appropriate metric for a given situation ensures that the acoustic behaviour can either be quantified after the fact to verify correct performance Ultraino or included in the solver to allow explicit optimisation of these quantities VolDisplaySingleParticle; Diff-PAT; AcousticElements. Once again, a versatile selection leads to more versatile applications for developers.
Method Solutions per Second
Analytical 4234
Finite Differences 1922
Table 2: Solutions per second for the two mentioned methods to compute Gor’kov Potential, an analytical solution that uses the derivative of the piston model to compute the gradient of pressure and the finite differences method which approximates this gradient. The analytical method is significantly faster over 3000 random independent points, computed sequentially.
@@
IV.5 Hardware
The final stage of the full-stack is Hardware, sending the transducer activations to a real world device. As discussed, this is difficult for a software package to be fully general purpose as every device may have a different communication protocol with little standardisation. Therefore, AcousTools uses OpenMPD’s board drivers OpenMPD to drive the 16x16 transducer arrays used in that paper. This is because these arrays are common across a number of applications, from levitationVolDisplaySingleParticle; GSPAT, haptics HapticsMPSTM; ToBEMorNot to fabricationAcoustoFab, have been the target of many phase retrieval methods Diff-PAT; GSPAT; VolDisplaySingleParticle and propagation methods AcousticElements; HighSpeedBEM, and are based on an open source platformUltraino; OpenMPD. Sending the activations is simple with AcousTools. All the developer must do is connect to the device and send the computed activations:
from acoustools.Levitator import
                    LevitatorController
from acoustools.Utilities import
                    create_points
from acoustools.Solvers import gspat
@@ IV.5 Hardware
IV.6 Application: Droplet-Solid Levitation
@@
The acoustic levitation of droplets is desirable from applications from fabricationLevitate; AcoustoFab to mid-air chemistryAcousticChemistry but is more challenging than solid particle levitation because too strong of a force applied to the droplet may cause it to buckle and burst Droplets; HighSpeedBEM. This can be mitigated by optimising the trap to have a reduced Gor’kov potential leading to controlled force to the droplet. On the other hand, the solid particle is best trapped by minimising Gor’kov potential and therefore applying high, converging forces. If traps of the same strength are used for both traps, the droplet would be at risk of bursting if both are strong or the solid bead would have sub-optimal forces if the traps are weaker, meaning if the particle is moved then it will be at risk of being droppedVolDisplaySingleParticle. Careful control of the forces gives a best-of-both-worlds situation, with the forces being as strong as possible for both traps without harming reliability Droplets.
A variety of different setups have been deployed for liquid levitation, from small single-axis devicesTinyLev; AcousticChemistry to small rings of transducers MultiFunctionalDroplets and arrays with solid reflectorsAcoustoFab; LeviPrint; HighSpeedBEM as well as multiple different propagatorsTinyLev; HighSpeedBEM and solvers MultiFunctionalDroplets; HighSpeedBEM being used across droplet levitation. Software that facilitates choices in these areas means that different contexts can be tested, compared and improved.
As has been discussed, AcousTools provides implementations that other software does not. Nether OpenMPD or Ultraino can optimise Gor’kov potential directly or simulate scattering and OpenMPD wouldn’t be capable of even defining some of the boards that find a home in droplet levitation.
AcousTools can achieve all the stages of the full-stack – from defining the context up to physicalising the levitation. By using an implicit-gradient descent solver to simply define and optimise an objective that computes the Gor’kov potential at two points, two different traps can be created. While this objective seeks to minimise the potential at one point, it minimises the square error between the potential at the other point and a given target value. Using an appropriate coupling parameter to keep both terms of a similar order of magnitude ensures that both terms are optimised with similar weight and produces two traps, one strong for the solid particle and the other with a weaker trapping strength for the droplet. When rendered in reality, the droplet stays intact, and both the particle and droplet can be rendered together. The objective for this can be described mathematically as:
@@
Here, $\bm{E}_{1}$ and $\bm{E}_{2}$ are the BEM propagators that model the scattering from a flat reflector to each of the target points. The resulting levitation as well as the simulated pressure and Gor’kov potential are shown in Fig. 7. This was conducted using a target potential of $U_{target}=-1\times 10^{-7}J$ for the droplet trap and a coupling parameter of $\lambda=7\times 10^{8}$ for consistent magnitudes between the two terms. A learning rate of $1$ was used and the optimisation was performed over $1000$ steps with the Adam optimiserAdam. The droplet trap has a resulting Gor’kov potential, $U=-9.98\times 10^{-8}J$, $99.8\%$ of the desired target.
@@
V Discussion
The acoustic full-stack gives a framework that researchers can use to evaluate their acoustic applications. When developing a novel method for some layer in the stack, it can be used to narrow down on comparable existing solutions. This framework will give clarity to field of acoustic holography and provide a lens to view it through. It will help new developers to know how to go about creating a new application, defining the points they need to decide between and, as adoption increases, helps them to quickly sort between what is applicable.
@@ V Discussion
Developers from a wide variety of fields can take AcousTools and deploy it in their contexts. Python as a language is among the easiest languages to learn pythonEase and is one of the most used languages PythonUse; StackOverflowSurvey meaning there is a very large number of people who would be able to take AcousTools and integrate acoustic holography into their own domain. A haptician could compare the effects of different transducer board or different propagators on immersion, someone interested in levitation could optimise the shapes that they render to give higher visual quality or a chemist could control small amounts of liquid drops for non-contact mid-air reactions while maintaining specific forces. AcousTools’ use of PyTorch not only provides the ability to integrate implicit-gradient solvers as discussed above but also could allow a user to deploy a machine learning solutions to their work. This means they can integrate novel solutions, some of which have already shown to be helpful for various acoustic domains AcousNet; StableLev. Thanks to AcousTools, possible applications in acoustic holography are not limited by existing software and now depend on developers alone.
We hope that AcousTools will become the standardised code library for acoustic holography for all applications, which does not currently exist. Using consistent code across researchers and industry developers would facilitate greater interoperability and reproducibility, which is currently hindered by the fact that developers and researchers mostly use their individual implementations. Of all existing software, AcousTools is the clear choice for this standard due to the aforementioned versatility and accessibility. This also means that as a greater degree of standardisation occurs, more work can be put into AcousTools by the whole community, extending the features it supports and making it even more suitable to be the standard software library. For example, one area that it may need extension is the support for other hardware, but the boards it does support are currently one of the most common hardware used OpenMPD, meaning this is only a minor drawback and is one shared with every other software package as well.
VI Conclusion
@@
Acknowledgements.
This work was supported by the UK Research and Innovation through their Frontier Research Guarantee Grant (No. EP/X019519/1), the Royal Academy of Engineering through their Chairs in Emerging Technology Program (No. CIET 17/18) and through the EPSRC prosperity partnership (No. EP/V037846/1).
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Joshua Mukherjee: Conceptualization (lead); Formal Analysis (lead); Investigation (lead); Methodology (lead); Software (lead); Validation (lead); Visualisation (lead); Writing - origional draft (lead); Writing - review & editing (lead). Giorgos Christopoulos: Writing - review & editing (supporting). Zhouyang Shen: Writing - review & editing (supporting). Sriram Subramanian: Conceptualization (supporting); Funding acquisition (equal); Supervision (supporting). Ryuji Hirayama: Conceptualization (supporting); Funding acquisition (equal); Supervision (lead). Writing - review & editing (supporting)
Data Availability Statement
AcousTools is open source and can be found at the following github repository: https://github.com/JoshuaMukherjee/AcousTools. The specific code used to generate figures from this paper can be found at the following github repository: https://github.com/JoshuaMukherjee/AcousTools_Paper

### v6

@@ 4.1 Setup
    = mesh_to_board(
        ’./data/Sphere-lam1.stl’)
This code loads the mesh and collects the positions and normals. In addition to custom transducer arrays, Acoustools has three common transducer configurations pre-configured, all based on arrays of 256 transducers arranged into 16x16 grids. This includes a single array positioned below the working volume and directed up [AcousticElements], a single array positioned above the working volume and pointed down [HighSpeedBEM], and one with the combination of these two arrays pointing towards one another [GSPAT]. These are supported as pre-constructed variables without needing to load a mesh, for example:
@@
After the boards are defined, points at which sound should be focused. To define a single point, $N=1$, located at $\boldsymbol{z}=(x_{t},y_{t},z_{t})$ can be simply defined as
p = create_points(N=1,x=xt,y=yt,z=zt)
@@
4.2 Propagators
As discussed above, the two main propagators used for acoustic holography are the piston model and the boundary element method. The piston model, as described in (1), uses a free-space model and is implemented in AcousTools as
F = forward_model_batched(p,board)
@@ 4.2 Propagators
Because these propagators share the same structure - both are matrices where the complex pressure at a point is obtained by the matrix-vector product of the propagator and the transducer activations - these can be swapped without requiring additional changes to the rest of the code. This allows a user to compare the two contexts with minimal changes to their code. For some solver (e.g., IB[IB]), the propagator is the only component that needs to be interchanged between the piston model and BEM contexts.
@@
4.3 Solvers
AcousTools provides a number of solvers, giving further flexibility depending on the context of an application. High-speed solvers, such as multipoint naive and GS-PAT, offer quick computation[GSPAT], while slower solvers, such as WGS, can generate focal points with small variance of pressure [TGS]. As discussed, the choice of which solver to use is a critical stage in the development process, and a wide variety that can be easily compared enables a developer to make an informed choice.
@@ 4.3 Solvers
AcousTools can leverage implicit-gradient descent algorithms to optimise more complex objectives [Diff-PAT, Diff-PAT-Expanded], which can go beyond what projective algorithms can do. Implicit gradient-descent methods can minimise Gor’kov potential directly, produce fields with specific metrics at specific points, and even combine multiple objectives simultaneously. Although these methods are often slower than others [Diff-PAT] and may not be suitable for all cases, their ability to use custom objectives provides greater flexibility for the developer.
@@
Then this can be sent to the implicit-gradient descent solver to be optimised, the exact optimiser and parameters can be customised such as the optimiser, the learning rate, weights of each term in the objective, and learning rate schedulers. These parameters can then be passed to the solver:
@@
        objective_params={
            ’lambda’:1.32e10},
@@
This results in the sound field shown in Fig. 4, using the Adam optimiser[Adam], a learning rate of $0.01$, 10001000 iterations and $\lambda=1.32\times 10^{10}$ which ensures each component has a similar magnitude. One point has a minima in the Gor’kov potential corresponding to a low pressure region sandwiched between two high pressure regions and the other point has a maxima of Gor’kov potential corresponding to a region of high pressure, as defined in the objective. This could be useful for multi-modal applications where a focal point is used for haptic applications while the minima of Gor’kov potential corresponds to a levitation trap that can hold a small bead [AcousticElements].
Projective solvers, such as IBP and GS-PAT, could not fully create the field in this situation, as they produce high pressure regions at their targets. These can be converted to traps through levitation signatures [AcousticElements], but such signatures convert all focal points instead of only one. As seen in Fig. 4, a focal point has a local minima just below the high pressure regions, and so a similar result to Fig. 4 could be produced by focusing slightly above the target trap position. However, using this method may lead to both sub-optimal solutions with smaller forces and is harder to customise the relative strengths of the trap and focus. Therefore, the implicit-gradient solver is an additional tool for a developer, and combining this with the other methods provides a full toolbox to use.
@@ 4.4 Analysis
pressure = propagate_abs(x, p, board=board)
U = gorkov(x, p, board=board)
@@
pressure = propagate(x, p, board=board)
Computing Gor’kov potential here uses analytical gradients of the piston model instead of using finite differences and makes it significantly faster compared to the standard method using finite differences [AcousticElements, HighSpeedBEM]. The speed-up between the two methods is significant, the analytical method is able to compute 4232 solutions per second compared to 1922 when using finite differences; these values were computed over 3000 random independent points, computed sequentially. A similar analytical solution is provided using the analytical gradient of BEM. For both propagators, an analytical method for computing the gradient of Gor’kov potential is also provided for computing force. A summary of these metrics can be seen in Fig. 5, showing that all these metrics from pressure to stiffness. In this case, the propagator used is the piston model and metrics use the analytical gradients where possible, similar visulisations could be created using BEM by simply swapping the propagators as discussed above.
For a developer, the ability to select the appropriate metric for a given situation ensures that the acoustic behaviour can either be quantified after the fact to verify correct performance [Ultraino] or included in the solver to allow explicit optimisation of these quantities [VolDisplaySingleParticle, Diff-PAT, AcousticElements]. Once again, a versatile selection leads to more versatile applications for developers.
@@
4.5 Hardware
The final stage of the full-stack is Hardware, sending the transducer activations to a real world device. As discussed, this is difficult for a software package to be fully general purpose as every device may have a different communication protocol with little standardisation. Therefore, AcousTools uses OpenMPD’s board drivers [OpenMPD] to drive the 16x16 transducer arrays used in that paper. This is because these arrays are common across a number of applications, from levitation[VolDisplaySingleParticle, GSPAT], haptics [HapticsMPSTM, ToBEMorNot] to fabrication[AcoustoFab], have been the target of many phase retrieval methods [Diff-PAT, GSPAT, VolDisplaySingleParticle] and propagation methods [AcousticElements, HighSpeedBEM], and are based on an open source platform[Ultraino, OpenMPD]. Sending the activations is simple with AcousTools. All the developer must do is connect to the device and send the computed activations:
@@ 4.5 Hardware
4.6 Application: Droplet-Solid Levitation
@@
The acoustic levitation of droplets is desirable from applications from fabrication[Levitate, AcoustoFab] to mid-air chemistry[AcousticChemistry] but is more challenging than solid particle levitation because too strong of a force applied to the droplet may cause it to buckle and burst [Droplets, HighSpeedBEM]. This can be mitigated by optimising the trap to have a reduced Gor’kov potential applying a controlled force to the droplet. On the other hand, the solid particle is best trapped by minimising Gor’kov potential and therefore applying high, converging forces. If traps of the same strength are used for both traps, the droplet would be at risk of bursting if both are strong or the solid bead would have sub-optimal forces if the traps are weaker, meaning if the particle is moved then it will be at risk of being dropped[VolDisplaySingleParticle]. Careful control of the forces gives a best-of-both-worlds situation, with the forces being as strong as possible for both traps without harming reliability [Droplets].
As AcousTools can achieve all the stages of the full-stack – from defining the context up to physicalising the levitation – droplet levitation can be easily deployed using an implicit-gradient descent solver. By defining an objective that computes the Gor’kov potential at two points, two different traps can be created. This objective seeks to minimise the potential at one point while it minimises the square error between the potential at the other and a given target value. Using an appropriate coupling parameter produces two traps, one strong for the solid particle and the other with a weaker trapping strength for the droplet. When rendered in reality, the droplet stays intact, and both the particle and droplet can be rendered together. The objective for this can be described mathematically as:
@@
Here, $\boldsymbol{E}_{1}$ and $\boldsymbol{E}_{2}$ are the BEM propagators that model the scattering from a flat reflector to each of the target points. The resulting levitation as well as the simulated pressure and Gor’kov potential are shown in Fig. 7. This was conducted using a $U_{target}=-1\times 10^{-7}J$ for the droplet trap and $\lambda=7\times 10^{8}$. A learning rate of $1$ was used and the optimisation was performed over $1000$ steps with the Adam optimiser[Adam]. The droplet trap has a resulting Gor’kov potential, $U=-9.98\times 10^{-8}J$, $99.8\%$ of the desired target.
@@
5 Discussion
The acoustic full-stack gives a framework that researchers can use to evaluate their acoustic applications. When developing a novel method for some layer in the stack, it can be used to narrow down on comparable existing solutions. This framework will give clarity to field of acoustic holography and provide a lens through which to view their work.
@@ 5 Discussion
Developers from a wide variety of fields can take AcousTools and deploy it in their contexts. Python as a language is among the easiest languages to learn [pythonEase] and is one of the most used languages [PythonUse] meaning there is a very large number of people who would be able to take AcousTools and integrate acoustic holography into their own domain. A haptician could compare the effects of different transducer board or different propagators on immersion, someone interested in levitation could optimise the shapes that they render to give higher visual quality or a chemist could control small amounts of liquid drops for non-contact mid-air reactions while maintaining specific forces. AcousTools’ use of PyTorch not only provides the ability to integrate implicit-gradient solvers as discussed above but also allows users to deploy a machine learning solutions to their work. Novel solutions that have already been shown to be helpful for various acoustic domains [AcousNet, StableLev]. Thanks to AcousTools, possible applications in acoustic holography are not limited by existing software and now depend on developers alone.
We hope that AcousTools will become the standardised code library for acoustic holography for all applications. Using consistent code across researchers and industry developers would facilitate greater interoperability and reproducibility, which is currently hindered by the fact that users mostly use their individual implementations. Of all existing software, AcousTools is the clear choice for this standard due to the aforementioned versatility and accessibility. As a greater degree of standardisation occurs, more work can be put into AcousTools by the whole community, extending the features it supports and making it even more suitable to be the standard software library.
6 Conclusion